1. Write log^2(2y=5)-1/2log^2(y=3) as a single log

2. find the exponential growth for a city whose population was 51,300 in 1980 and 240,500 in 2000

Answer by stanbon(75887)

1. Write log^2(2y=5)-1/2log^2(y=3) as a single log
-----
log2[(2y+5)/(y+3)^(1/2)]
------------------------------
2. find the exponential growth for a city whose population was 51,300 in 1980 and 240,500 in 2000
----
You have 2 points: (0, 51,300) and (20, 240,500)
Form: y = ab^x
Using (0,51,300) : 51,300 = ab^0
a = 51,300
------
Using (20,240,500) : 240,500 = 51,300b^20
b^20 = 4.688
---
b = 4.688^(1/20)
b = 1.08
--------
Equation:
y = 51,300*1.08^x
-------------------------
Cheers,
Stan H.
=========================